# Shortcut to Chemically Accurate Quantum Computing via Density-based Basis-set Correction

### v1

@@ III Conclusion
We demonstrated the applicability of the DBBSC method to quantum computing algorithms for chemistry. Using ADAPT-VQE, these approaches were shown to be able to systematically improve STO-3G minimal basis-set results by predicting total energies that are intermediate between double-zeta and triple-zeta FCI qualities. Overall, the presented self-consistent total energies are fully in line with their a posteriori counterparts, opening therefore a path towards computations of properties thanks to the availability of improved QC densities. As the a posteriori approach can be virtually applied to any QC computation performed on real hardware via classical computing, it provides an affordable improvement strategy of current QC computations. Triple-zeta accuracy can be achieved using double-zeta basis sets or by using our SABS approach. Such “black-box” pivoted Cholesky strategy for the on-the fly generation of basis sets has been shown to be competitive and often superior to available standard choices while offering systematic computational savings over large basis sets reducing therefore significantly the qubit requirements. Their usefulness is therefore not limited to quantum computing algorithms as they also offer systematically improvable solutions for performing reference classical computations. Indeed, since SABS are user-defined truncated versions of the Dunning basis sets that can match a qubit budget, they offer a reduced cost access to improved accuracy for any quantum chemistry method. Overall, the DBBSC method allows for a partition of the Hilbert space treatment which can remain handled by the QC ansatz embedded in an effective DFT correction dealing with the missing short-range electron correlation effects while the missing HF basis-set correction energy is obtained via a computationally affordable process. In such a framework, the two parts of the basis-set correction do not require to be evaluated at the QC level and can be fully shifted to classical computing while retaining the core QC algorithm features at a constant qubit count. The strategy allows us to access an improved electronic density thanks to a self-consistent scheme opening a path for the evaluation of first-order properties such as dipole moments. Of course, since the approach is hybridized with DFT, it is focused on improving the density and does not provide the exact many-body wave function. However, it provides an accelerated basis-set convergence, strongly improves the solutions and allows us to compute chemically meaningful energies and properties on systems that would have required far more than 100 logical qubits. We also verified that the correction gets smaller and smaller as the basis set increases so it has the correct CBS limit. Consequently, using a DBBSC/SABS approach, the computation of the H2 total energy at the FCI/CBS level that would have required more than 220 logical qubits, i.e. the qubit count needed to perform a least a FCI/cc-pV5Z computation, can be achieved here with only 24 qubits. Overall, we were able to converge four systems to the FCI/CBS limit including He, Be, H2, and LiH. We were also able to provide accurate dissociation curves for H2, LiH, and N2. Such computations required the use of only a single GPU. Since most quantum chemistry studies are presently out of reach of quantum computers, this research opens the path to more affordable quantitative quantum chemistry simulations of small molecules using QC algorithms. Thus, a posteriori corrections can be virtually added to any type of STO-3G VQE fermionic computations on real hardware [5, 48, 49] allowing us to improve significantly their accuracy at very little computational cost. Since adaptive simulations on real hardware are making some progress [50] while hardware itself improves, high-resolution corrected simulations should be progressively possible on future quantum computers providing a route to FCI quality computations. This strategy is particularly suited for resources demanding computations that converge slowly and require large basis sets associated to large qubits counts. Finally, DBBSC approaches are not limited to ground-state computations and can be extended to excited states via a linear response formalism[15]. This will be the subject of future research in the context of quantum computing. In that context, beside the state-of-the-art ADAPT-VQE hybrid quantum-classical algorithm, it would be interesting to revisit accurate fixed-ansätze [51], such as UCCSDT [52] (and others), to analyze their convergence when coupled to the DBBSC method. In practice, the presented DFT basis-set correction framework is not restricted to a given QC ansatz and can leverage any future QC algorithmic improvements. Thus, our present DBBSC/SABS methodology still requires to use qubits either through quantum hardware or through the use of a classical quantum emulator. The qubit count is therefore our main limitation. The computations of this paper used up to 32 qubits and represent a proof-of-concept study of what one can do with quantum emulation. However, state-vector simulations have limitations due to memory. Indeed, if they are theoretically possible up to 40-50 qubits on very large exascale supercomputers, they start to become relatively unpractical for chemical simulations when reaching 36 qubits due to the high computational resources and time-to-solution requirements. To explore further the chemical space with quantum algorithms, we are currently upgrading our Hyperion-1 framework to increase our emulated qubit counts by going beyond the state-vector formalism thanks to various elements from the CUDA-Q SDK [53, 54] developed by NVIDIA. We are also presently testing various implementations of the DBBSC algorithms on available quantum hardware. To conclude, by reducing the number of qubits required to reach the CBS limit, we expect to tackle predictive real-world quantum chemistry applications with strategies applicable to both NISQ and FTQC quantum computing algorithms.
Acknowledgment

### v2

@@ III Conclusion
We demonstrated the applicability of the DBBSC method to QC algorithms for quantum chemistry. Using ADAPT-VQE, these approaches were shown to be able to systematically improve minimal basis-set results by predicting ground-state energies that are intermediate between double-zeta and triple-zeta FCI qualities. Overall, the presented self-consistent basis-set corrected ground-state energies are very close to their non-self-consistent counterparts. As the latter a posteriori basis-set correction approach can be very easily applied to any wave-function QC calculation performed on real quantum hardware, it provides an affordable improvement strategy for current QC chemistry computations. The self-consistent basis-set correction scheme is still useful since it permits the calculation of properties such as dipole moments thanks to the availability of improved QC densities.
An additional reduction of the required basis size is provided by our SABS approach. Such a “black-box” pivoted-Cholesky strategy for the on-the fly generation of basis sets has been shown to be competitive and often superior to available standard choices. It offers systematic computational savings for large basis sets, reducing significantly the qubit requirements. Thus the SABS’ usefulness is not limited to QC algorithms as they also offer systematically improvable solutions for performing reference classical computations. Indeed, since SABS are user-defined truncated versions of the Dunning basis sets that can match a qubit budget, they also offer a reduced cost access to improved accuracy for any quantum-chemistry method.
Our methodology allows us to compute chemically meaningful energies and properties on systems that would have required far more than 100 logical qubits. For example, the computation of the H2 total energy at the FCI/cc-pV5Z level that would have required more than 220 logical qubits, can be achieved here with only 24 qubits using our basis-set correction scheme and our SABS technique. Overall, we were able to converge four systems to the FCI/CBS limit including He, Be, H2, and LiH. We were also able to provide accurate dissociation curves for H2, LiH, and N2. Such computations required the use of only a single GPU. Since most quantum-chemistry studies are presently out of reach of quantum computers, this research opens the path to more affordable quantitative quantum-chemistry simulations of small molecules using QC algorithms. In particular, the present a posteriori basis-set corrections can be easily added to any type of STO-3G VQE fermionic computations on real hardware [5, 53, 54], allowing one to improve significantly their accuracy at very little computational cost. Since adaptive simulations on real hardware are making some progress [55] while hardware itself improves, basis-set corrected simulations should be progressively possible on future quantum computers providing a route to FCI/CBS quality computations. This strategy is particularly suited for resources demanding computations that converge slowly and require large basis sets associated to large qubits counts. Finally, the DBBSC method is not limited to ground-state computations and can be extended to excited states via a linear-response formalism[17]. This will be the subject of future QC research.
@@
In this context, beside the state-of-the-art ADAPT-VQE hybrid quantum-classical algorithm, it would be interesting to revisit accurate fixed wave-function ansätze [56], such as UCCSDT [57] and others, to analyze their convergence when coupled to the DBBSC method. In practice, the presented basis-set correction framework is not restricted to a given QC wave-function ansatz and can leverage any future QC algorithmic improvements. Our present DBBSC/SABS methodology still requires to use qubits either through quantum hardware or through the use of a classical quantum emulator. The qubit count is therefore our main limitation. The computations of this paper used up to 32 qubits and represent a proof-of-concept study of what one can do with quantum emulation. However, state-vector simulations have limitations due to memory. Indeed, if they are theoretically possible up to 40-50 qubits on very large exascale supercomputers, they start to become relatively unpractical for quantum-chemistry simulations when reaching 36 qubits due to the high computational resources and time-to-solution requirements. To explore further the chemical electronic space with quantum algorithms, we are currently upgrading our Hyperion-1 framework to increase our emulated qubit counts by going beyond the state-vector formalism thanks to various elements from the CUDA-Q SDK [58, 59] developed by NVIDIA. We are also presently testing various implementations of the DBBSC algorithms on available quantum hardware. To conclude, by reducing the number of qubits required to reach the CBS limit, we expect to tackle predictive real-world quantum-chemistry applications with strategies applicable to both NISQ and FTQC algorithms.
Acknowledgments